# Binding Affinity Prediction: From Conventional to Machine Learning-Based Approaches

### v1

@@ 2.2 Binding Affinity Datasets
PDSP (Psychoactive Drug Screening Program) Ki Database. The PDSP Ki Database (Roth et al., 2000), founded in 2006, serves as a regularly updated data warehouse for published binding affinity data of drugs and drug candidates for receptors, ion channels, transporters, and enzymes. The entries are accessible for query under given conditions or can be downloaded as a CSV file. For each entry, the IDs of the ligand and receptors are provided, along with additional information such as the SMILES string and the reference for the affinity data. Despite the large quantity of affinity data, there are two major drawbacks to the PDSP Ki Database. First, the affinity data ($K_{i}$ value) for each compound-receptor pair may be either a single numeric value or an inequality such as “greater than 7000.” Secondly, structure features for both protein and ligand are not included in the dataset, meaning users will have to query other data sources such as PDB.
@@
KIBA (Kinase Inhibitor Bioactivity) Dataset was originally proposed in Tang et al. (2014). It aimed to integrate various bioactivity types, including $K_{i}$, $K_{d}$, and $IC_{50}$, into a single term named the KIBA score, and to demonstrate its usage in classifying kinase inhibitor targets and pinpointing errors in binding affinity datasets. The dataset was obtained from searching for drug-kinase interactions in ChEMBL with $IC_{50}$ and at least one binding constant terms ($K_{i}$ and $K_{d}$). For each interaction in the KIBA dataset, they have converted the bioactivity terms into KIBA score using the equations below:
@@
KIBA =$=\begin{cases}{K_{i}}_{\cdot adj}&\text{if ${IC}_{50}$ and $K_{i}$ are present}\\ {K_{d}}_{\cdot adj}&\text{if ${IC}_{50}$ and $K_{d}$ are present}\\ ({K_{i}}_{\cdot adj}+{K_{d}}_{\cdot adj})/2&\text{if ${IC}_{50}$, $K_{i}$, and$K_{d}$ are present}\end{cases}$ (2)
@@
with $L_{i}=0.3$ and $L_{d}=1.3$, which are the optimal values for adjustment. There are no features of any kind included in the original dataset. Users need to look up ChEMBL, PDB, or other sources to obtain features of drug-kinase pairs.
@@
Other Binding Affinity Datasets. AffinDB (Block et al., 2006) was a major binding affinity dataset when it was proposed in 2006. However, the website and affinity data have not been updated for some time. As of 2019, there are 748 affinity values covering 474 PDB complexes, which is severely insufficient compared to PDBbind or BindingDB. CSAR (Community Structure-Activity Resource) (Smith et al., 2011; Dunbar et al., 2011; Damm-Ganamet et al., 2013; Dunbar et al., 2013; Smith et al., 2016; Carlson et al., 2016) was an experimental dataset of crystal structures and binding affinities for diverse protein-ligand complexes. The dataset was intended for a community-wide exercise conducted by a group at the University of Michigan from 2010 to 2014, aiming to use unpublished data from in-house projects to evaluate existing protein-ligand binding affinity prediction methods. The CSAR dataset has inspired some of the most meaningful discussions on the methods and evaluations of binding affinity, such as those in (Novikov et al., 2011; Koes et al., 2013; Carlson, 2016), and the ideas and insights from these discussions are still relevant today. Aside from those datasets that are no longer maintained, there are many datasets that are relevant to binding affinity in one aspect or another. For instance, Tang et al. (2014) cite studies of bioactivity profiling of small-molecule protein kinase inhibitor by Davis et al. (2011), Metz et al. (2011), and Anastassiadis et al. (2011), each of which provides a set of measurements of binding affinity between kinase and small compounds. Krivák and Hoksza (2018) combined multiple small protein-ligand complex datasets, such as CHEN11 (Chen et al., 2011), COACH420 (Roy et al., 2012; Yang et al., 2013b), and HOLO4K (Schmidtke et al., 2010), for training and evaluation of binding site prediction. As existing datasets and models fail to account for the dynamic features of protein-ligand interactions, the recently released PLAS-20K (2024) addresses this gap. As an extension of PLAS-5K (2022), PLAS-20K includes 97,500 independent simulations across 19,500 different protein-ligand complexes.
@@ 3 Methodology
Next, we discuss the various protein-ligand binding affinity prediction methodologies as developed from the early 2000s to the present day. We divided all the methods into two categories based on the model complexity and input data. In roughly chronological order, they are: (1) Conventional methods, which are essentially a set of equations derived from assumptions and understanding of the binding process, combining weighted physio-chemical terms in an additive manner into a single estimate for affinity. (2) Machine learning models (e.g., random forest, support vector machine, neural network, etc.) trained with human-engineered descriptors extracted from the protein-ligand complexes; and representation-learning methods, which extract features directly from raw data of protein-ligand complexes (e.g., SMILES strings, voxels, graphs) using the learning capacity of deep neural networks. Roughly speaking, as the study of binding affinity advances, the community has shifted from methods that require extensive domain knowledge and assumptions to those that can exploit the increasing number of available protein-ligand complex structures.
@@ 3.1 Conventional Methods
Ever since the formalization of the concept of protein-ligand docking, researchers have been trying to predict the binding affinity or energy based on human understanding of physics and chemistry. These methods were often called scoring functions, rather than predictors or models, because they almost always tool the forms of additive functions, combining various engineered physio-chemical terms. Traditionally, these conventional scoring functions have been roughly divided into the following three categories, in multiple works (Cheng et al., 2009; Ballester and Mitchell, 2010a; Ashtawy and Mahapatra, 2012):
@@ 3.1.1 Physics-based methods
Traditionally, these methods are often referred as force-field-based methods, because “classical force-field-based scoring functions assess the binding energy by calculating the sum of the non-bonded interactions” (Meng et al., 2011), which is often done with force field. However, Liu and Wang (2015) suggested the change of name from force-field methods to physics-based methods because some of the methods or models commonly used in this category of scoring functions, such as Poisson-Boltzmann or Generalized Born continuum solvation models, end-point approximation, and quantum mechanics computation, are not based on force fields. The new name, physics-based methods, more appropriately reflects the theoretical basis of this approach. In sum, physics-based methods calculate the binding energy by adding up the Van der Waals, electrostatic, hydrogen bonding (which could be included implicitly in electrostatic terms), and desolvation energy terms (Liu and Wang, 2015). The general functional form of physics-based methods is shown in the following equation:
$\Delta G_{\text{binding}}=\Delta E_{\text{vdw}}+\Delta E_{\text{electrostatic}}+(\Delta E_{\text{H-bond}})+\Delta G_{\text{desolvation}}$ (3)
Physics-based methods take advantage of their ab initio nature, where no additional information about the protein and ligand is required other than the binding pose. The interactions can be described using molecular force fields or quantum chemistry calculations.
Physics-based methods for binding affinity prediction consider detailed atomic interactions and capture subtle effects like solvation, entropic contributions, and conformational flexibility. These methods offer mechanistic insights into binding processes, identifying key residues and interactions critical for binding. They can also account for complex cases involving significant ligand and protein conformational changes. When accurately parameterized, these methods can reliably provide energy landscapes for novel compounds, helping to understand the kinetics and thermodynamics of binding (Kuhn et al., 2020; Guterres and Im, 2020; Fu et al., 2022).
However, these methods are computationally intensive and time-consuming, often requiring substantial computational resources and long simulation times (Guterres and Im, 2020). Simulation results are also sensitive to the choice of force fields, simulation parameters, and initial conditions (Fu et al., 2022). Their application is often limited to small to medium-sized systems due to the high computational cost. Additionally, achieving convergence in simulations can be very difficult, which can affect the reliability of binding affinity predictions (Fu et al., 2022).
@@
Empirical methods are the most commonly used score-based models, found in packages such as Rosetta and Autodock (Rohl et al., 2004; Trott and Olson, 2010). In addition to the binding energy terms from the previous physics-based methods, they include contributions from empirical chemistry factors, such as hydrophobicity, metal-ligand interactions, entropy effects from steric hindrance, and ligand motifs (Rohl et al., 2004; O’Meara et al., 2015; Park et al., 2016; Alford et al., 2017; Halgren et al., 2004; Friesner et al., 2006; Quiroga and Villarreal, 2016). These methods employ statistical learning algorithms such as multivariate linear regression (MLR) or partial least squares regression (PLS) to parameterize the individual energy terms with weights to estimate the binding affinity. Generally speaking, empirical methods are more flexible and adaptive, allowing users to add customized terms to the scoring function. They adopt a generalized functional form such as:
$Score=E_{bind}+E_{emp}$ (4)
where $E_{bind}$ is binding affinity from the physics model and $E_{emp}$ represents empirical terms that account for interactions and contributions that pure physics fails to capture due to the limitations of force field parameters. The latter additional terms improve the binding affinity prediction based on pure physics models and can be tuned to achieve a better fit for particular interests, such as pharmaceutical importance (Wang et al., 2004b; Friesner et al., 2006). However, it requires empirical knowledge to set up the right scoring functions and training to optimize the weight of each term.
@@
As the name suggests, knowledge-based potential methods rely on learning from a database of protein-ligand complexes to determine the potential between atom pairs and predict binding affinity. The main assumption behind these methods is that if an atom pair, one from the ligand and one from the protein, appears with a higher frequency than the reference distribution, it might indicate an energetically favorable interaction between the given pair. Thus, the distance-dependent potential of an atom pair can be obtained through inverse Boltzmann analysis based on the measured occurrence frequency of this pair across the entire knowledge base.
In some works, these methods are also referred to as knowledge-based methods (Gohlke et al., 2000) or potential of mean force methods (Su et al., 2009). The general functional form of knowledge-based potential methods is represented by the following equation:
@@
where $k_{B}$ is the Boltzmann constant, $T$ is the absolute temperature, and $r$ is the distance between atom pairs. The knowledge-based methods are more general, implicitly incorporating effects not fully understood from the structural data. Additionally, they can also incorporate some of the physics-based or empirical terms to enhance performance.
Knowledge-based potential methods have the the following advantages: (1) They are computationally simple. (2) To obtain the distributions of atom-pair occurrences, these methods can take advantage of complex structures without measured binding affinity. Their disadvantages are as follows: (1) The distribution of atom pairs is only relevant to the binding pose rather than binding affinity, which theoretically makes knowledge-based potential methods inferior to other methods. However, in practice, this drawback does not seem to hurt performance, as knowledge-based potential methods are comparable to others. (2) The distribution statistics obtained from the knowledge base are not accurate and should not be assumed to be in a Boltzmann distribution. This means that the potential terms in the functional forms are only approximations.
3.1.4 Hybridization of these three methods

### v2

@@ 2.2 Binding Affinity Datasets
PDSP (Psychoactive Drug Screening Program) Ki Database. The PDSP Ki Database (Roth et al., 2000), founded in 2006, serves as a regularly updated data warehouse for published binding affinity data of drugs and drug candidates for receptors, ion channels, transporters, and enzymes. The entries are accessible for query under given conditions or can be downloaded as a CSV file. For each entry, the IDs of the ligand and receptors are provided, along with additional information, such as the SMILES string and the reference for the affinity data. Despite the large quantity of affinity data, there are two major drawbacks to the PDSP Ki Database. First, the affinity data ($K_{i}$ value) for each compound-receptor pair may be either a single numeric value or an inequality, such as “greater than 7000.” Secondly, structure features for both protein and ligand are not included in the dataset, meaning users will have to query other data sources such as PDB.
@@
KIBA (Kinase Inhibitor Bioactivity) Dataset. KIBA was originally proposed in Tang et al. (2014). It aimed to integrate various bioactivity types, including $K_{i}$, $K_{d}$, and $IC_{50}$, into a single term named the KIBA score, and to demonstrate its usage in classifying kinase inhibitor targets and pinpointing errors in binding affinity datasets. The dataset was obtained from searching for drug-kinase interactions in ChEMBL with $IC_{50}$ and at least one binding constant terms ($K_{i}$ and $K_{d}$). For each interaction in the KIBA dataset, they have converted the bioactivity terms into KIBA score using the equations below:
@@
[…]
@@
with $L_{i}=0.3$ and $L_{d}=1.3$, which are the optimal values for adjustment. There are no features included; users need to look up ChEMBL, PDB, or other sources to obtain features of drug-kinase pairs.
@@
Other Binding Affinity Datasets. AffinDB (Block et al., 2006) was a major binding affinity dataset when it was proposed in 2006. However, the website and affinity data have not been updated for some time. As of 2019, there are 748 affinity values covering 474 PDB complexes, which is significantly fewer than PDBbind and BindingDB. CSAR (Community Structure-Activity Resource) (Smith et al., 2011; Dunbar et al., 2011; Damm-Ganamet et al., 2013; Dunbar et al., 2013; Smith et al., 2016; Carlson et al., 2016) was an experimental dataset of crystal structures and binding affinities for diverse protein-ligand complexes. The dataset was intended for a community-wide exercise conducted by a group at the University of Michigan from 2010 to 2014, aiming to use unpublished data from in-house projects to evaluate existing protein-ligand binding affinity prediction methods. CSAR has inspired some of the most meaningful discussions on the methods and evaluations of binding affinity (Novikov et al., 2011; Koes et al., 2013; Carlson, 2016), which remain relevant today. Aside from these unmaintained datasets, there are many datasets that are relevant to binding affinity. For instance, Tang et al. (2014) cite studies of bioactivity profiling of small-molecule protein kinase inhibitor by Davis et al. (2011), Metz et al. (2011), and Anastassiadis et al. (2011), each of which provides a set of measurements of binding affinity between kinase and small compounds. Krivák and Hoksza (2018) combined multiple small protein-ligand complex datasets, such as CHEN11 (Chen et al., 2011), COACH420 (Roy et al., 2012; Yang et al., 2013b), and HOLO4K (Schmidtke et al., 2010), for training and evaluation of binding site prediction. CrossDocked2020 (Francoeur et al., 2020), which contains 22.5 million ligand poses docked into diverse but structurally related binding pockets from the Protein Data Bank, is also widely employed as a benchmark dataset for evaluating protein–ligand binding affinity prediction methods. As existing datasets and models fail to account for the dynamic features of protein-ligand interactions, the recently released PLAS-20K (2024) addresses this gap. As an extension of PLAS-5K (2022), PLAS-20K includes 97,500 independent simulations across 19,500 different protein-ligand complexes. Nevertheless, PLAS-20K remains one of the few publicly available molecular dynamics resources, and its scale is still modest compared to static datasets such as PDBbind. Most other MD-based datasets are either proprietary or highly heterogeneous, limiting their utility as community benchmarks. This scarcity of large, standardized dynamic datasets represents a key bottleneck, as models trained only on static snapshots often fail to capture critical factors such as protein flexibility, induced fit, and entropic effects.
While PDBbind has been the de facto benchmark dataset for binding affinity prediction, it is well recognized that it suffers from significant issues of data leakage and redundancy, which limit the generalizability of models trained on it. This concern becomes particularly evident when evaluating models on proprietary or out-of-distribution industrial datasets, where performance often drops substantially. Recent efforts have emphasized the need for more robust benchmarks that explicitly address data leakage and bias. For instance, initiatives such as Plinder and discussions in the community (Open Molecular Software Foundation, 2023) have highlighted directions for curating cleaner, more diverse datasets with stricter train–test splits to improve real-world applicability. Moreover, some traditional resources like Binding MOAD are no longer actively maintained (Hu and Lill, 2023), further underscoring the importance of developing next-generation benchmark datasets. Future progress in binding affinity prediction will thus depend not only on improved modeling techniques but also on the availability of high-quality datasets designed to avoid redundancy and leakage.
@@ 3 Methodology
Next, we discuss the various protein-ligand binding affinity prediction methodologies developed from the early 2000s to the present day. We divided all the methods into two categories based on model complexity and input data. (1) Conventional methods, which are essentially a set of energy equations derived from assumptions and understanding of the binding process, combining weighted physio-chemical terms in an additive manner into a single estimate for affinity. (2) Machine learning models (e.g., random forest, support vector machine, neural network) trained with human-engineered descriptors extracted from the protein-ligand complexes; and representation-learning methods, which extract features directly from raw data of protein-ligand complexes (e.g., SMILES strings, voxels, graphs) using the learning capacity of deep neural networks. As the study of binding affinity advances, the community has shifted from methods that require extensive domain knowledge and assumptions to those that can exploit the increasing number of available protein-ligand complex structures.
@@ 3.1 Conventional Methods
Since the formalization of the concept of protein-ligand docking, researchers have been trying to predict the binding affinity or energy based on human understanding of physics and chemistry. These methods were often called scoring functions, rather than predictors or models, because they almost always tool the forms of additive functions, combining various engineered physio-chemical terms. Traditionally, these conventional scoring functions have been roughly divided into the following three categories (Cheng et al., 2009; Ballester and Mitchell, 2010a; Ashtawy and Mahapatra, 2012): Physics-based methods, empirical methods, and knowledge-based potential methods. Among these, physics-based methods represent the most direct attempt to calculate binding affinity from first principles, and thus form a natural starting point for our discussion.
@@ 3.1.1 Physics-based methods
Traditional calculations of protein–ligand binding affinity that are grounded in statistical mechanics can be divided into two main categories. Both ultimately connect to the standard definition of the binding free energy,
$\Delta G_{b}^{\circ}=-k_{\mathrm{B}}T\ln\!\big(C^{\circ}K_{b}\big),$ (3)
where $K_{b}$ is the equilibrium association constant and $C^{\circ}$ is the 1 M standard concentration.
The first category is commonly referred to as free energy surface (FES) or potential of mean force (PMF) approaches. These methods compute a reduced-dimensional free energy profile along selected reaction coordinates that capture the essential physics of the binding process. The resulting profile reveals minimum free-energy pathways and critical points such as intermediates and transition states, from which both equilibrium and kinetic information can be extracted. In the simplest radial case, the binding constant can be obtained from the PMF as
$K_{b}=\int_{\text{site}}4\pi r^{2}\,e^{-\beta[\,w(r)-w(r^{\ast})\,]}\,dr,$ (4)
where $w(r)$ is the PMF and $r^{\ast}$ is a reference position in bulk solvent.
The second category is alchemical free energy (AFE) methods (Ngo et al., 2024). Here, the focus is not on the real physical pathway but on exploiting the fact that free energy is a state function. The ligand is gradually “alchemically” decoupled from its surroundings, both in the binding site and in bulk solvent, and the difference provides the binding free energy. A general expression can be written as
$\Delta G_{b}^{\circ}=\big[\Delta G^{\text{site}}_{\text{int}}-\Delta G^{\text{bulk}}_{\text{int}}\big]+\Delta G_{\text{restraints}},$ (5)
where the first term accounts for the difference in interaction free energies between site and bulk, and $\Delta G_{\text{restraints}}$ includes standard-state, translational, rotational, and conformational corrections. Alchemical methods are often more computationally efficient and robust for estimating $\Delta G_{b}^{\circ}$, but they sacrifice mechanistic insight into the actual binding pathway.
Relative binding free energy (RBFE) methods. While the alchemical approach refers to absolute binding free energy (ABFE) calculations, in practice the relative binding free energy method has become far more widely adopted, particularly in industrial drug discovery settings. Rather than computing the binding affinity of a single ligand from scratch, RBFE calculates the free energy difference between two structurally similar ligands binding to the same target:
$\Delta\Delta G=\Delta G_{\text{b}}^{\text{ligand B}}-\Delta G_{\text{b}}^{\text{ligand A}}.$ (6)
This is achieved by performing alchemical transformations that mutate ligand A into ligand B, both in the binding site and in solvent. The relative formulation dramatically reduces the sampling burden: because structurally similar ligands often occupy overlapping conformational spaces and induce similar protein reorganization, many systematic errors and convergence challenges cancel out in the difference. This makes RBFE substantially more computationally tractable than ABFE for comparing congeneric series of compounds (Wang et al., 2015; Cournia et al., 2017).
As a result, RBFE remains the gold standard for prospective ligand optimization in pharmaceutical settings, with demonstrated success in predicting relative potencies with chemical accuracy (± 1 kcal/mol) (Behera et al., 2025).
In practice, however, both FES- and AFE-based approaches are limited by the inherently low sampling efficiency of molecular dynamics simulations (York, 2023). As binding often involves slow conformational changes and rare transitions, straightforward Boltzmann sampling cannot ensure sufficient exploration of the relevant configurational space.
To overcome this challenge, these methods are commonly combined with enhanced sampling techniques, such as replica exchange (Sugita and Okamoto, 1999), umbrella sampling (Torrie and Valleau, 1977), or metadynamics (Laio and Parrinello, 2002), which improve convergence by facilitating transitions across free energy barriers and ensuring adequate sampling of binding-relevant degrees of freedom.
However, exhaustive sampling of all relevant conformational states is often prohibitively expensive for realistic biomolecular systems. This limitation has motivated the development of so-called endpoint methods, which bypass the explicit sampling of the binding pathway by considering only the end states: the bound complex and the separated receptor and ligand. Among these, the MM/PBSA and MM/GBSA (Kollman et al., 2000) approaches have become particularly popular, owing to their balance between computational cost and predictive accuracy. In these methods, the binding free energy is expressed as
$\Delta G_{\mathrm{bind}}=\Delta E_{\mathrm{MM}}+\Delta G_{\mathrm{solvation}}-T\Delta S,$ (7)
where $\Delta E_{\mathrm{MM}}$ contains bonded, electrostatic, and van der Waals terms, $\Delta G_{\mathrm{solvation}}$ is decomposed into polar and non-polar contributions, and $T\Delta S$ represents the entropic term.
The distinction between MM/PBSA and MM/GBSA lies in the treatment of the polar solvation energy. In MM/PBSA, it is obtained by numerically solving the Poisson–Boltzmann (PB) equation, whereas MM/GBSA employs the Generalized Born (GB) approximation:
$\Delta G_{\mathrm{solvation}}=G_{\mathrm{pol}}^{\mathrm{PB/GB}}+G_{\mathrm{np}}^{\mathrm{SASA}}.$ (8)
Here, the polar solvation free energy, $G_{\mathrm{pol}}^{PB/GB}$, is obtained either by numerically solving the Poisson–Boltzmann (PB) equation or by employing the Generalized Born (GB) approximation. In the PB formulation, the electrostatic potential $\phi(\mathbf{r})$ is obtained from
$\nabla\cdot\left[\epsilon(\mathbf{r})\nabla\phi(\mathbf{r})\right]+\kappa^{2}(\mathbf{r})\sinh\!\left(\phi(\mathbf{r})\right)=4\pi\rho(\mathbf{r}),$ (9)
and the corresponding solvation energy is calculated as
$G_{\mathrm{pol}}^{PB}=\frac{1}{2}\sum_{i}q_{i}\,\phi(\mathbf{r}_{i}).$ (10)
In the GB approximation, the polar solvation energy is estimated analytically as
$G_{\mathrm{pol}}^{GB}=-\frac{1}{2}\left(1-\frac{1}{\epsilon}\right)\sum_{i=1}^{N}\sum_{j=1}^{N}\frac{q_{i}q_{j}}{f_{ij}},$ (11)
with
$f_{ij}=\sqrt{r_{ij}^{2}+\alpha_{i}\alpha_{j}\exp\!\left(-\frac{r_{ij}^{2}}{4\alpha_{i}\alpha_{j}}\right)},$ (12)
where $r_{ij}$ is the interatomic distance, and $\alpha_{i}$ are the effective Born radii.
The principal advantages of these endpoint approaches are their efficiency and modularity: they require only molecular mechanics snapshots, can be combined with standard MD simulations, and allow energy decomposition analysis. Nevertheless, their accuracy is strongly system-dependent, with limitations arising from approximate entropy estimates, neglect of conformational reorganization and water thermodynamics, and sensitivity to continuum solvent parameters. Thus, while MM/PBSA and MM/GBSA represent an attractive compromise between speed and accuracy, their predictive power remains limited compared to fully converged alchemical free energy methods. While physics-based methods strive to compute binding free energies directly from first principles, their high computational cost and dependence on approximations have led to the development of empirical methods, which instead introduce experimentally motivated terms and statistical regression to improve efficiency and practical accuracy.
@@
Empirical methods are the most commonly used score-based models, found in packages such as Rosetta and Autodock (Rohl et al., 2004; Trott and Olson, 2010). In addition to the binding energy terms from the previous physics-based methods, they include contributions from empirical chemistry factors, such as hydrophobicity, metal-ligand interactions, entropy effects from steric hindrance, and ligand motifs (Rohl et al., 2004; O’Meara et al., 2015; Park et al., 2016; Alford et al., 2017; Halgren et al., 2004; Friesner et al., 2006; Quiroga and Villarreal, 2016).
These methods employ statistical learning algorithms, such as multivariate linear regression (MLR) or partial least squares regression (PLS), to parameterize the individual energy terms with weights for estimating the binding affinity. Generally, empirical methods are more flexible and adaptive, allowing users to add customized terms to the scoring function. They adopt a generalized functional form:
$\mathrm{Score}=E_{\mathrm{bind}}+E_{\mathrm{emp}}$ (13)
where $E_{bind}$ is binding affinity from the physics model and $E_{emp}$ represents empirical terms that account for interactions and contributions that pure physics fails to capture due to the limitations of force field parameters.
Hydrophobic Interaction Terms
Hydrophobic interactions constitute one of the most critical empirical components. AutoDock Vina implements hydrophobic terms using a distance-dependent step function (Trott and Olson, 2010):
$E_{\mathrm{hydrophobic}}=w_{\mathrm{hydrophobic}}\sum_{i,j}f_{\mathrm{hydrophobic}}(r_{ij},t_{i},t_{j})$ (14)
Metal-Ligand Interaction Terms
Metal coordination represents a specialized empirical term crucial for metalloproteins. The general form incorporates both distance and angular constraints (Santos-Martins et al., 2014):
$E_{\mathrm{metal}}=\sum_{M,L}w_{\mathrm{ML}}\cdot g(r_{\mathrm{ML}},r_{\mathrm{opt}},\sigma)\cdot h(\theta,\theta_{\mathrm{ideal}})$ (15)
where the distance function follows a Gaussian form:
$g(r_{\mathrm{ML}},r_{\mathrm{opt}},\sigma)=\exp\left(-\frac{(r_{\mathrm{ML}}-r_{\mathrm{opt}})^{2}}{2\sigma^{2}}\right)$ (16)
and the angular function penalizes deviations from ideal coordination geometry:
$h(\theta,\theta_{\mathrm{ideal}})=\exp\left(-\alpha(\theta-\theta_{\mathrm{ideal}})^{2}\right)$ (17)
Entropy Effect Terms
Empirical entropy terms account for the loss of conformational and rotational freedom upon binding. The rotational entropy loss is approximated using the rigid rotor model (Gilson et al., 1997):
$\Delta S_{\mathrm{rot}}=-R\ln\left(\frac{8\pi^{2}I_{A}I_{B}I_{C}}{\sigma_{\mathrm{rot}}h^{3}}\right)$ (18)
where $I_{A}$, $I_{B}$, $I_{C}$ are the principal moments of inertia, $\sigma_{rot}$ is the rotational symmetry number, and $h$ is Planck’s constant.
For conformational entropy, AutoDock Vina employs a simplified model based on the number of rotatable bonds (Trott and Olson, 2010):
$E_{\mathrm{entropy}}=w_{\mathrm{tors}}\cdot\frac{N_{\mathrm{tors}}}{1+w_{\mathrm{tors}}\cdot N_{\mathrm{tors}}}$ (19)
where $N_{tors}$ is the number of rotatable bonds and $w_{tors}$ represents the entropic penalty per rotatable bond.
The latter additional terms improve the binding affinity prediction based on pure physics models and can be tuned to achieve a better fit for particular interests, such as pharmaceutical importance (Wang et al., 2004b; Friesner et al., 2006). However, empirical methods require substantial empirical knowledge to set up appropriate scoring functions and extensive training datasets to optimize the weight of each term. The challenge lies in balancing model complexity with generalizability, as overfitting to training data can compromise performance on novel chemical scaffolds or binding sites (Li et al., 2014b). While empirical methods enhance physics-based scoring by incorporating experimentally motivated terms and statistical regression, they still require substantial expert knowledge to define and parameterize interaction terms. To reduce this dependence on manual feature design, researchers developed knowledge-based potential methods that instead derive interaction preferences directly from large databases of protein–ligand complexes.
@@
Knowledge-based potential methods rely on learning from a database of protein-ligand complexes to determine the potential between atom pairs and predict binding affinity. The main assumption behind these methods is that if an atom pair, one from the ligand and one from the protein, appears with a higher frequency than the reference distribution, it might indicate an energetically favorable interaction between the given pair. Thus, the distance-dependent potential of an atom pair can be obtained through inverse Boltzmann analysis based on the measured occurrence frequency of this pair across the entire knowledge base. These methods are sometimes referred to as knowledge-based methods (Gohlke et al., 2000) or potential of mean force methods (Su et al., 2009). The general functional form of knowledge-based potential methods is represented by:
@@
where $k_{B}$ is the Boltzmann constant, $T$ is the absolute temperature, and $r$ is the distance between pairs of atoms. Knowledge-based methods are more general, implicitly incorporating effects not fully understood from the structural data. They can also incorporate some physics-based or empirical terms to enhance performance.
Knowledge-based potential methods offer several distinct advantages:
(1) Computational efficiency: These methods are computationally lightweight and fast, making them suitable for high-throughput screening applications where rapid evaluation of large compound libraries is essential.
However, these methods face significant limitations:
(1) Structural bias over energetic relevance: The statistical distributions of atom-pair distances are primarily reflective of geometrically favorable binding poses rather than true thermodynamic binding affinity. This theoretical limitation suggests that knowledge-based potentials may not capture the energetic determinants of binding strength. Paradoxically, empirical evaluations demonstrate that these methods often perform comparably to more sophisticated approaches, indicating that structural complementarity may be a reasonable proxy for binding affinity in many cases.
(2) Limited physical interpretability: While knowledge-based potentials represent a significant advancement by introducing data-driven elements to complement purely physics-based or empirical approaches, they remain fundamentally phenomenological. The statistical potentials lack clear physical meaning and cannot easily be related to specific intermolecular forces or thermodynamic properties. Recognition of these complementary strengths and limitations has motivated the development of hybrid scoring functions that strategically combine knowledge-based potentials with physics-based calculations and empirical observations, aiming to harness the computational efficiency of statistical methods while maintaining physical rigor and predictive accuracy.
3.1.4 Hybrid methods